### v1

# Wenbin Wang, Wenjie Xu, Colin N. Jones

Abstract
Human-aware controllers play an important role in engineering systems for improving productivity, efficiency, and sustainability. It is essential to design such a controller that optimizes user utility while adhering to plant dynamics. While most online optimization algorithms rely on first-order or zeroth-order oracles, human feedback often appears as pairwise comparisons. In this work, we propose an online feedback optimization algorithm that leverages such preference feedback. We design a controller that estimates the gradient based on the binary pairwise comparison result between two consecutive points and study its coupled behavior with a nonlinear plant. Under mild assumptions on both the utility and the plant dynamics, we establish explicit stability criteria and quantify sub-optimality. The theoretical findings are further supported through numerical experiments.
I Introduction
People are the key components in engineering systems and are the primary beneficiaries of many leading technologies, such as smart building energy management systems [1], human-robot-collaboration [2]. Designing a human-aware controller capable of regulating the system in real time to optimize user latent utility is essential. Existing controller typically tracks a predefined reference, which is often derived from large population models, e.g., indoor room temperature [3]. While simple and easy to implement, it can introduce bias and operate the system at a suboptimal point, as it fails to account for individual differences. Moreover, without real-time human feedback, such a controller cannot adapt to time-varying utility and is not robust to external disturbances.
As an emerging real-time control technique, Online Feedback Optimization (OFO) [4] has proven effective in applications such as power system operation [5] and robot coordination [6]. By taking real-time system output, OFO can navigate the system toward its optimal point without requiring precise knowledge of plant dynamics or measurement noise. Theoretical guarantees have been established for both first-order [7] and zeroth-order [8] oracles. However, people are generally more adept at making relative comparisons than providing absolute evaluations of their utility [9]. As a result, human feedback often appears as pairwise comparisons, making it difficult to directly apply existing OFO schemes to the design of a human-aware controller.
Research on offline optimization with preference feedback has been rapidly expanding. In the case of a finite action space, human preference can be represented by a preference matrix where each entry indicates the probability of one option being preferred over another. Various definitions of optimal point have been proposed, e.g., Copeland winner [10] and Borda winner [11], with methods like random exploration [12] and greedy search [13] employed to identify the optimal choice. For continuous action spaces, a common assumption is that the preference is drawn from a latent utility function, which can be modeled via Gaussian Processes (GP). Several heuristic strategies for sequential decision-making have been developed to balance exploration and exploitation [14, 15], while regret guarantees can be established by considering functions within a Reproducing Kernel Hilbert Space (RKHS) [16, 17]. Additionally, various gradient estimation techniques leveraging preference feedback have been explored [18, 19], with optimality guarantees under regularity assumptions such as smoothness and convexity [19].
Despite recent advancements, many open challenges remain in developing online preference optimization algorithms that account for plant transients. First, the stability of the interconnected system between the controller and the plant is hard to quantify. This is not only due to the binary nature of preference feedback, which makes the system highly nonlinear, but also due to the fact that people have limited knowledge about the plant. Directly following their preferences can make the system unstable. Moreover, a human utility function can be time-varying and environment-dependent [9]. With external disturbances, it is difficult for a controller trained on offline data to track the optimal point in real time. The way different options are presented also has an impact on how people express their preferences given that people are not always rational [9]. A carefully designed mechanism for human-controller interaction is necessary to ensure efficient system operation.
Inspired by the recently proposed model-free OFO with one-point residual estimation method [8], we introduce a novel gradient estimator based on preference feedback. Our approach employs a randomized scheme, where a random exploration signal is added into the plant at each time step. The preference feedback is then collected from the perturbed utility function between the current and the previous time steps. Different from existing work [18, 19] where two function evaluations are required at each time step, our method requires one evaluation, making it well-suited for online implementation. Under some mild assumptions, we show that the resulting update imitates a gradient descent step with respect to a probability function and provide theoretical guarantees for the closed-loop system. The theoretical results are further validated through numerical experiments on a thermal comfort optimization problem.
II Problem Formulation and Preliminaries
II-A Notation
Let $\mathbb{R}^{n}$ be the $n$-dimensional Euclidean space. For a real-valued function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, we denote its gradient at $x\in\mathbb{R}^{n}$ by $\nabla f(x)$. For a single input function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$, we denote its derivative by $\sigma^{\prime}(t)$. We denote the value of vector $x\in\mathbb{R}^{n}$ at time step $k$ by $x_{k}$.
II-B Problem Formulation
Consider an exponentially stable plant with $n_{x}\in\mathbb{N}$ states and $n_{u}\in\mathbb{N}$ inputs
$x_{k+1}=f(x_{k},u_{k}),$ (1)
where $f:\mathbb{R}^{n_{x}\times n_{u}}\rightarrow\mathbb{R}^{n_{x}}$ is the state transition function, $x\in\mathbb{R}^{n_{x}}$ is the plant state, $u_{k}\in\mathbb{R}^{n_{u}}$ is the plant input. We assume that there exists a unique steady-state input-state map $h:\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{x}}$ such that $\forall u\in\mathbb{R}^{n_{u}}$, $h(u)=f(h(u),u)$.
Assumption 1.
The input-state map $h$ is $L_{h}$-Lipschitz continuous.
The assumption on the exponentially stable plant with Lipschitz continuous input-state map is satisfied for many systems, e.g., linear system $x_{k+1}=Ax_{k}+Bu_{k}$ with $A$ being invertible and stable. The result of this work can also be extended to include the output measurement $y_{k}=g(x_{k},u_{k})$, where additional assumptions such as Lipschitz continuous input-output map are necessary.
According to the converse Lyapunov theorem [8, 20, P. 194], for an exponentially stable plant, there exists a Lyapunov function $V:\mathbb{R}^{n_{x}\times n_{u}}\rightarrow\mathbb{R}$ and positive parameters $\alpha_{1},\alpha_{2},\alpha_{3}$ such that $\forall u\in\mathbb{R}^{n_{u}}$,
$\alpha_{1}\|x-h(u)\|^{2}\leq V(x,u)\leq\alpha_{2}\|x-h(u)\|^{2},$ (2)
$V(f(x,u),u)-V(x,u)\leq-\alpha_{3}\|x-h(u)\|^{2}.$ (3)
The parameter $\alpha_{3}$ characterizes how fast the system stabilizes. A larger $\alpha_{3}$ indicates the system stabilizes to $h(u)$ with a faster rate.
In this work, we want to optimize user utility when the plant is at the steady state
$\begin{split}\min_{x,u}\quad&\Phi(x,u),\\ \text{s.t.}\quad&x=h(u),\end{split}$ (4)
where $\Phi:\mathbb{R}^{n_{x}\times n_{u}}\rightarrow\mathbb{R}$ is the latent utility function. This utility function depends on both the system state and input, capturing real-world scenarios where these factors jointly have influences. For example, in building energy management, there is an inherent trade-off between thermal comfort and control effort [1].
Assumption 2.
$\forall u\in\mathbb{R}^{n_{u}}$, the latent utility $\Phi(x,u)$ is $L_{x}$-Lipschitz with respect to the first argument, i.e., $\|\Phi(x_{1},u)-\Phi(x_{2},u)\|\leq L_{x}\|x_{1}-x_{2}\|,\forall x_{1},x_{2}\in\mathbb{R}^{n_{x}}$, where $L_{x}$ is a uniform Lipschitz constant.
We can reformulate (4) as an unconstrained problem by replacing $x$ with $x=h(u)$ in $\Phi(x,u)$ and get
$\min_{u}\tilde{\Phi}(u),$ (5)
where $\tilde{\Phi}(u)=\Phi(h(u),u)$.
Assumption 3.
The differentiable function $\tilde{\Phi}(u)$ is $L_{0}$-Lipschitz, $L_{1}$-smooth and $m$-strongly convex.
Assumption 3 is commonly employed in the literature both for theoretical analysis [21, 22, 23] and applications, e.g., building control [1] and frequency regulation [24]. We denote the solution of (5) by $u^{*}$. The assumption on strongly convexity assures that $u^{*}$ is unique.
Classical numerical optimization methods, such as gradient descent, require the exact knowledge of $h$ to find the optimal solution [4]. However, this is impractical for systems where constructing an accurate physical model is costly, such as buildings. To address this, zeroth-order feedback optimization methods have been developed to solve (4). These methods estimate the gradient using finite differences, either considering transients [8] or assuming an algebraic plant [22]. However, directly evaluating $\Phi(x,u)$ from human feedback at a given state and input can introduce bias [25], violating the common assumption of noiseless objective evaluation in zeroth-order optimization.
The goal in this work is to develop an effective online algorithm that can regulate the system (1) towards $u^{*}$ with preference feedback. Different from the existing setting, preference feedback often appears in the form of pairwise comparison, a binary signal. For the problem to be properly defined, we need to link such binary feedback with the utility through a user model.
II-C User model
Given $u_{1}$ and $u_{2}$, we denote the event ‘$\tilde{\Phi}(u_{1})<\tilde{\Phi}(u_{2})$’ by $u_{1}\succ u_{2}$, i.e., $u_{1}$ is better than $u_{2}$. Preference feedback is denoted $\bm{1}_{u_{1}\succ u_{2}}$, which takes the value
$\bm{1}_{u_{1}\succ u_{2}}=\begin{cases}1,&\text{if $u_{1}$ is better},\\ -1,&\text{if $u_{2}$ is better}.\end{cases}$ (6)
To link the utility $\tilde{\Phi}$ with preference feedback, we adopt a probabilistic model as shown in Assumption 4.
Assumption 4.
Preference feedback $\bm{1}_{u_{1}\succ u_{2}}$ follows a Bernoulli distribution, i.e., $\mathbb{P}(\bm{1}_{u_{1}\succ u_{2}}=1)=\sigma(\tilde{\Phi}(u_{2})-\tilde{\Phi}(u_{1}))$, where $\sigma(t)=\frac{1}{1+e^{-t}}$.
This model, i.e., Bradley-Terry Model, among other probabilistic models such as Thurstone-Mosteller model [26, Section 2.2.3], are commonly found in the literature for preferential learning [18, 16]. The idea is that when $\tilde{\Phi}(u_{1})$ is smaller, the probability of reporting $\bm{1}_{u_{1}\succ u_{2}}=1$ is larger. We denote the Lipschitz constant and smoothness constant of $\sigma(t)$ by $L_{\sigma,0}$, $L_{\sigma,1}$, respectively. Furthermore, the result in this work also holds with other linking functions $\sigma(t)$, provided they are monotonically increasing, satisfy $\sigma(-\infty)=0$, $\sigma(\infty)=1$, are rotation-symmetric, and exhibit convexity for $x\leq 0$, essentially behaving like cumulative distribution functions [18].
III Online gradient estimation with preference feedback
We build on the existing result in model-free feedback optimization [8] and design a controller that establishes a gradient estimator with preference feedback from real-time state measurement. The binary preference feedback is collected from the user between two consecutive function evaluations. The controller then regulates the system toward the point where the utility is minimized. This is summarized in Algorithm 1
Algorithm 1 Preferential learning algorithm
0:  step size $\eta$, smoothing parameter $\delta$, $u_{0}\in\mathbb{R}^{n_{u}}$, number of time steps $T$
1:  for $k=1,\ldots,T-1$ do
2:     $x_{k+1}=f(x_{k},u_{k}+\delta v_{k})$;
3:     ask user to express their preference between $\Phi(x_{k+1},u_{k}+\delta v_{k})$ and $\Phi(x_{k},u_{k-1}+\delta v_{k-1})$ and sample $\bm{1}_{(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})}$;
4:     update the input via $u_{k+1}=u_{k}+\frac{\eta}{2\delta}\bm{1}_{(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})}v_{k}$;
5:  end for
5:  $u_{T}$
At time step $k$, a random exploration signal $v_{k}$ is added to the plant as shown in Line 2 of Algorithm 1. It is drawn from a $(n_{u}-1)$-dimension unit sphere $\mathbb{S}^{n_{u}-1}$. We then use $x_{k+1}$ as the approximation of $h(u_{k}+\delta v_{k})$, which are under the influence of $v_{k}$. This allow us to approximate $\tilde{\Phi}(u_{k}+\delta v_{k})$ with $\Phi(x_{k+1},u_{k}+\delta v_{k})$.
Then in Line 3, user is asked to provide preference over the pairs, $\Phi(x_{k+1},u_{k}+\delta v_{k})$ and $\Phi(x_{k},u_{k-1}+\delta v_{k-1})$, i.e., sampling the random variable $\bm{1}_{(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})}$. We slightly abuse notation here, denoting the event ‘$\Phi(x_{k+1},u_{k}+\delta v_{k})<\Phi(x_{k},u_{k-1}+\delta v_{k-1})$’ by $(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})$ and representing the preference feedback via random variable $\bm{1}_{(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})}$, where $\mathbb{P}(\bm{1}_{(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})}=1)=\sigma(\Phi(x_{k},u_{k-1}+\delta v_{k-1})-\Phi(x_{k+1},u_{k}+\delta v_{k}))$.
Finally, the control input is updated as shown in Line 4. When $\Phi(x_{k+1},u_{k}+\delta v_{k})$ is lower, $\mathbb{P}(\bm{1}_{(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})}=1)$ is higher, which means with high probability the algorithm will set $u_{k+1}$ closer to $u_{k}+\delta v_{k}$, where the utility is lower.
The resulting closed-loop system between the plant and the controller takes the form
$x_{k+1}=$ $f(x_{k},u_{k}+\delta v_{k}),$ (7)
$u_{k+1}=$ $u_{k}+\frac{\eta}{2\delta}\bm{1}_{(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})}v_{k}.$ (8)
The preference feedback reveals the descent direction of the unknown utility. Randomizing the strategy by adding $v_{k}$ to the system facilitates exploration of all possible directions at a given point, effectively imitating a gradient descent step. Although (8) is not estimating $\nabla\tilde{\Phi}(u)$ since preference feedback does not contain information on actual function difference, we will demonstrate that the algorithm performs gradient descent on a probability function, ultimately seeking the point with the highest winning probability.
IV Performance Analysis
We first show that with controller (8), system (7) is stable.
Lemma 1.
Let Assumptions [1-4] hold. $\mathbb{E}[V(x_{k},u_{k}+\delta v_{k})]\leq\mu^{k}\mathbb{E}[V(x_{0},u_{0}+\delta v_{0})]+\frac{a_{1}}{1-\mu}(2\delta^{2}+\eta+(\frac{\eta}{2\delta})^{2})$, where $\mu=\frac{2\alpha_{2}}{\alpha_{1}}(1-\frac{\alpha_{3}}{\alpha_{2}})$, and $a_{1}=4\alpha_{2}L_{h}^{2}$.
Proof.
Proof can be found in Appendix -A. ∎
Remark 1.
We select the expected Lyapunov function $\mathbb{E}[V(x_{k},u_{k}+\delta v_{k})]$ as the indicator for stability. The decay rate $\mu$ is a function of $\alpha_{1},\alpha_{2},\alpha_{3}$, which represents how fast the system (1) stabilizes with a constant input. A smaller $\mu$ indicates the system will more rapidly stabilize to the equilibrium point. From Lemma 1, $\mathbb{E}[V(x_{k},u_{k}+\delta v_{k})]$ is upper bounded by an exponentially decaying component and a constant term. This is because the increment of $\|u_{k}\|$ is bounded by $\frac{\eta}{2\delta}$, which is summable for an exponentially stable plant. Consequently, the system cannot grow unbounded.
To simplify the notation, we compactly express the probability when $\tilde{\Phi}(u)$ is available for measure as $\mathbb{P}(\bm{1}_{u^{\prime}\succ u}=1)=p_{u^{\prime}}(u)=\sigma(\tilde{\Phi}(u)-\tilde{\Phi}(u^{\prime}))$. A smaller $\tilde{\Phi}(u)$ results in a lower $p_{u^{\prime}}(u)$. This monotonic relationship suggests that, rather than minimizing $\tilde{\Phi}(u)$ directly, one can first minimize $p_{u^{\prime}}(u)$, which is further supported by Lemma 2.
Lemma 2.
Let Assumptions [1-4] hold. $\forall u^{\prime}\in\mathbb{R}^{n_{u}}$, $p_{u^{\prime}}(u)$ is $L_{p,0}$-Lipschitz, and $L_{p,1}$-smooth with respect to $u$, where $L_{p,0}=L_{\sigma,0}L_{0}$, $L_{p,1}=\sigma^{\prime}(0)L_{1}+L_{\sigma,1}L_{0}^{2}$. Furthermore, $p_{u^{\prime}}(u)$ is partially-convex with respect to $u$ if $\tilde{\Phi}(u)\leq\tilde{\Phi}(u^{\prime})$.
Proof.
Proof on partially-convexity can be found in [18]. The rest proof can be found in Appendix -B. ∎
Since $\sigma(t)$ is a bounded smooth function, the Lipschitz and smoothness property is preserved for the composed function $p_{u^{\prime}}(u)$. Convexity is also partially preserved since $\sigma(t)$ is partially convex. Under the assumption on strongly convexity, we show that the minimizer of $p_{u^{\prime}}(u)$ for any fixed $u^{\prime}$ is unique and coincides with $u^{*}$.
Lemma 3.
Let Assumptions [1-4] hold and $u^{*}$ be the unique solution of (5), then $\forall u^{\prime}\in\mathbb{R}^{n_{u}}$, $u^{*}$ is also the unique minimizer of $p_{u^{\prime}}(u)$.
Proof.
Proof can be found in Appendix -C. ∎
Based on the conclusion from Lemma 2, 3 and techniques from convex analysis, we can prove that performing gradient descent with respect to $p_{u^{\prime}}(u)$
$u_{k+1}=u_{k}-\eta\nabla p_{u_{k}}(u_{k})$ (9)
will converge to $u^{*}$, where $\nabla p_{u_{k}}(u_{k})$ is the gradient of $p_{u_{k}}(u)$ with respect to $u$ evaluated at $u_{k}$. Indeed, the update (9) is the ideal case where knowledge of $h$ is available. It can be shown that $\nabla\tilde{\Phi}(u_{k})$ and $\nabla p_{u_{k}}(u_{k})$ are the same up to a constant factor. Building on this evidence, we compactly write (8) as
$u_{k+1}=u_{k}-\eta(\nabla p_{u_{k}}(u_{k})+e_{k}),$ (10)
where the error $e_{k}$ is defined as
$e_{k}=-\frac{1}{2\delta}\bm{1}_{(x_{k+1},u_{k}+\delta v_{k})\succ(x_{k},u_{k-1}+\delta v_{k-1})}v_{k}-\nabla p_{u_{k}}(u_{k}).$ (11)
In (10), we interpret (8) as an instance of (9) with an error term $e_{k}$ in gradient evaluation. The error exists since both $h(u_{k}+\delta v_{k})$ is approximated with $x_{k+1}$ and the gradient is estimated via random perturbation. We further show that $e_{k}$ is bounded in Lemma 4.
Lemma 4.
Let Assumptions [1-4] hold. $\|\mathbb{E}[e_{k}|\mathcal{F}_{k}]\|\leq\sqrt{R_{1}V(x_{k-1},u_{k-1}+\delta v_{k-1})+R_{2}}$, where $R_{1}=\frac{2L_{\sigma,0}^{2}L_{x}^{2}(\mu+1)}{\alpha_{2}\delta^{2}}\mu$, and $R_{2}=\frac{2L_{\sigma,0}^{2}L_{x}^{2}a_{1}}{\alpha_{2}\delta^{2}}(2\delta^{2}+\eta+(\frac{\eta}{2\delta})^{2})\mu+2a_{2}^{2}\delta^{2}$, $a_{2}=L_{p,1}\sqrt{n}+(\sigma^{\prime}(0)L_{1}+L_{p,1})(1+\frac{\eta}{\delta})$
Proof.
Proof can be found in Appendix -D. ∎
In Lemma 4, we observe that at time step $k$, conditioned on the natural filtration $\mathcal{F}_{k}$, the error $e_{k}$ is bounded by $R_{1}V(x_{k-1},u_{k-1}+\delta v_{k-1})$ and $R_{2}$, where $V(x_{k-1},u_{k-1}+\delta v_{k-1})$ is bounded as shown in (1). Additionally, $R_{1}=\mathcal{O}(\mu)$ and $R_{2}=\mathcal{O}(\mu,\delta^{2})$, both of which decrease as $\mu$ decreases. This result aligns with the approximation in Section III, as a smaller $\mu$ allows $x_{k+1}$ to better approximate $h(u_{k}+\delta v_{k})$.
Now we are ready to present the main stability result.
Theorem 1.
$\forall k^{\prime}$ and $\forall k>k^{\prime}$, the expected distance to $u^{*}$ is bounded by $\mathbb{E}[\|u_{k}-u^{*}\|^{2}]\leq(\frac{1+\rho}{2})^{k-k^{\prime}}\mathbb{E}[\|u_{k^{\prime}}-u^{*}\|^{2}]+\mathcal{O}(\mu,\mu^{k^{\prime}},\delta)$, where $\rho=1-2\sigma^{\prime}(0)m\eta$.
Proof.
Proof can be found in Appendix -E. ∎
We use the expected distance to $u^{*}$, i.e., $\mathbb{E}[\|u_{k}-u^{*}\|^{2}]$, as the error metric, as it is commonly employed in the literature for online zeroth-order optimization [22, 27]. Our results indicate that for a fixed time step $k^{\prime}$, $\mathbb{E}[\|u_{k}-u^{*}\|^{2}]$ is bounded by $\mathbb{E}[\|u_{k^{\prime}}-u^{*}\|^{2}]$, scaled by an exponentially decaying factor, along with a constant term $\mathcal{O}(\mu,\mu^{k^{\prime}},\delta)$, which is the function of $k^{\prime}$. At equilibrium, $\mathbb{E}[\|u_{k^{\prime}}-u^{*}\|^{2}]$ vanishes, leaving the left-hand side bounded solely by $\mathcal{O}(\mu,\mu^{k^{\prime}},\delta)$. Since $\mu^{k^{\prime}}$ decreases as $k^{\prime}$ increases for $\mu<1$, the steady-state error is eventually fully characterized by $\mathcal{O}(\mu,\delta)$.
V Numerical simulation
To show that the controller (8) is capable of identifying $u^{*}$, we conduct several simulations with an LTI system,
$\begin{split}x_{k+1}=Ax_{k}+Bu_{k},\end{split}$ (12)
where $A\in\mathbb{R}^{n_{x}\times n_{x}}$, $B\in\mathbb{R}^{n_{x}\times n_{u}}$ are the system matrices. This system has a steady-state input-output map $H=(I-A)^{-1}B$ and is pre-stabilized by a lower-level controller, ensuring that the spectral radius of $A$ is less than one. Such systems are commonly found in various applications, including building control [28], and power systems [24].
V-A A simple example
To analyze the impact of plant transient on algorithm performance, we first consider a quadratic problem
$\min$ $\quad(x-x_{\text{ref}})^{\top}(x-x_{\text{ref}}),$
$s.t.$ $\quad x=Hu,$
where the solution can be written as $u=H^{-1}x_{\text{ref}}$. We assign $A=\begin{bmatrix}c&1\\ 0&c\end{bmatrix}$, $B$ to be the identity matrix and $x_{\text{ref}}=[100,100]^{\top}$. We switch $c$ from $c=0.1$ to $c=0.7$ to represent plants with different decay rates. We set $\eta=0.1$ and $\delta=0.5$ in the simulation. Note that a quadric function does not have a bounded gradient. Hence, Assumption 2 is not satisfied. The result can be found in Fig. 1.
(a) c = 0.1.
(b) c = 0.7.
Figure 1: Quadratic problem.
In Fig. 1, the relative error $\|x_{k}-x_{\text{ref}}\|/\|x_{\text{ref}}\|$ in logarithmic scale is plotted for $c=0.1$ and $c=0.7$. The solid line represents the mean value over 20 simulations, while the transparent area is the standard deviation. The orange line represents the algebraic plant, i.e., assuming $H$ is known and directly sampling $\bm{1}_{u_{k}+\delta v_{k}\succ u_{k-1}+\delta v_{k-1}}$, and the blue line represents (8). When at the steady state, both (8) and its steady-state counterpart return the result with the same level of precision. However, for a slower system ($c\!=\!0.7$), (8) has a larger overshoot and a higher variance in steady state.
V-B Thermal comfort optimization
Next, we consider a thermal comfort optimization problem where a building is modeled via an LTI system with 13 states [28]. We model an occupant’s utility on thermal comfort with the well-known Predictive Mean Vote (PMV) model [3]. The PMV model is usually quantified via Predicted Percentage of Dissatisfied (PPD), which is a nonconvex function with respect to room temperature. We generate the preference feedback based on this model, aiming to explore different room temperatures to minimize PPD using controller (8). For other inputs required by the PMV model, we assume the occupant is typing and wearing sweatpants, T-shirt and Shoes or sandals (default setting in the PMV model).
In Fig. 2, the mean indoor temperature (blue solid line) is plotted over 20 runs of simulations. The black horizontal line represents the ground truth. With careful tuning of $\eta$ and $\delta$, we found that (8) can track the optimal point effectively without large overshoot, which demonstrates its potential in learning a human’s utility in real-world applications.
Figure 2: Thermal comfort optimization.
VI Conclusion
In this work, we designed a human-aware controller that utilizes preference feedback to optimize utility in a dynamic system. We derived an explicit upper bound on the error introduced by approximating the steady-state input-state map using real-state measurements and random perturbations. We analyzed the impact of both the system decay rate and smoothing parameters on the stability and optimality of the closed-loop system. Numerical experiments on thermal comfort optimization demonstrate its potential for solving real-world problems. Further directions include extending the theoretical framework to alternative user models, e.g., noise-free, and exploring real-world applications in areas like product design and chemical selection.